# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 4.1.1 Classical Anonymization
As already mentioned, anonymity definitions can be traced back to 1997 with k-anonymity. Anonymization aims to make impossible the re-identification of an individual, that is the process of reconstructing the identity of an individual given a record that contains their information. Different attributes could be anonymized. While effectively masking PII – like names, emails, and social security numbers – is necessary, some features that are not exactly identifiers can be used in combination to identify an individual: those features are called quasi-identifiers. Quasi-identifiers are apparently non-personal attributes – like gender, date of birth, and zip codes – that can be used in combination with external resources to identify a person from its anonymized records. The definition of k-anonymity (Sweeney, 1997) ensures that quasi-identifier cannot be used for re-identification. Essentially, a dataset is said to have k-anonymity if any given record in the dataset is indistinguishable from at least k - 1 other records with respect to certain identifying attributes called quasi-identifiers. Upon this concept, several implementation techniques were proposed by Samarati & Sweeney (1998), like generalization (modifying attribute values to less specific but accurate representations), suppression (i.e., entirely removing data that are outliers or too identifiable), minimal Generalization (ensuring data are not generalized more than necessary to meet k-anonymity). Of course, this method is highly susceptible to the choice of k (even inapplicable sometimes) and gives coarse data, making it impossible to carry out fine-grained analysis. It is also important to mention that, due to the reliance on quasi-identifiers, this method is much less effective with unstructured text.
@@
The standard process is outlined in Figure 4. After any necessary preprocessing phase, the first challenge to solve is to identify the entities that may cause a violation of privacy. This can be done via a Named Entity Recognition (NER) tool. To maintain the coherence of the text, a Co-reference Resolution (CRR) system can be used to resolve references between entities should be identified if the anonymization framework aims to ensure consistent anonymization across the document. Finally, private information should be replaced with anonymous information. The anonymization methods can vary. Suppression is used to completely remove the information. Tagging allows the replacement of sensitive information with artificial labels –like the one obtained by the NER system– that still retain information about the class of the data and an identifier. The sensitive information can be also replaced with other textual information: random substitution and generalization can both be used. Random substitution uses another random entity of the same class, while generalization replaces an entity with a more general term.
@@
More recent anonymization methods also leverage Deep Neural Networks. Zhou et al. (2022) use BERT to impute previously masked sensitive information. This framework is called DataSifterText and it works with a whitelist that holds sensitive terms that need to be masked and a blacklist for meaningful terms that should be retained. After a proportion parameter is chosen, sensitive tokens are masked according to that proportion, and then BERT is used to replace the masked tokens. These two steps in the procedure respectively cover the identification and anonymization step shown in 4).
@@
In some cases, the author of a text is meant to be identified: re-identification of an author can be done by spotting some features that characterize an author’s writing. Romanov & Fedotova (2020) analyze this scenario for the Russian language. They extract features like lemmas, parts of speech and morphological features. Then, a fast correlation filter – a multivariate feature selection method where the class relevance and the dependency between each feature pair are taken into account – is used to select the most informative features. Those informative features are the more likely to be identifying features. To preserve the privacy of the author, those features are smoothed by making targeted adjustments: for example, words or punctuation marks that are characteristic of the author’s style are replaced or altered based on average statistics derived from a broader corpus. The final step is to further process the smoothed text with a deep learning model: in this work, a Transformer-based model is chosen. This approach reduces the probability of authorship identification to random guessing.
@@
Mosallanezhad et al. (2019) use an attention-based task-aware text representation learner for extracting latent embedding representations with a reinforcement-learning-based privacy and utility preserver for manipulating these embeddings. While it effectively balances privacy and data utility and is adaptable and efficient, it has notable drawbacks, including complexity, potential data-specific limitations, and dependency on precise parameter tuning.
@@
More recently, some works proposed the use of LLMs for anonymization. Vats et al. (2023) propose to hide sensitive tokens via a masking process and impute them by using a Large Language model (LLM). In relation to 4 the masking acts as the identification part, as it finds and masks sensitive information, while the imputation represents the proper anonymization, where non-sensitive terms replace the (previously masked) sensitive ones. To carry out the masking process, several techniques are proposed: a handcrafted list of safe words (the ones that do not carry sensitive information) called allowlist created by linguistic experts, this is similar to the blacklist of Zhou et al. (2022) and identifies the terms that do not need to be masked; another method to identify safe words is vocabThres, where the N more frequent words in a vocabulary are considered safe; the last masking criteria is to find the unsafe words and involves using a NER module that identify potentially private tokens to mask them (similar to Mamede et al. (2016)). After the choice of a masking method, the chosen tokens are masked and then imputed by an LLM. Also for the imputation, three separate methods are proposed: Top-1, which uses the one token with the highest probability from the LLM’s output; Top-K, which randomly selects one of the first K tokens with the highest output probabilities; Fine-Tuning, which selects the records that do not contain sensitive information and uses them to fine-tune the LLM, then just takes the highest probability output (so after fine-tuning it circles back to the Top-1 approach). They compare models trained on the anonymized datasets against models trained on the masked dataset, showing a significant improvement in both tasks.
@@
As the usage of closed models increases, privacy concerns also emerge when sending prompts containing private information. Chen et al. (2023b) focus on sanitizing prompts before sending them out to some unsafe third party like a black-box model like ChatGPT, thus they enforce prompt privacy. Moreover, they argue that standard anonymization techniques – like masking – may be useful to ensure the privacy of prompts, but the model output could not be effectively de-anonymized to improve its utility. They propose a new framework called HaS (Hide and Seek) with two local small LMs, one trained for anonymization and one trained for deanonymization. The first one preprocesses the prompt before sending it to the black-box and the second one postprocesses the model output. The hide model is trained with original sentences and extracted privacy entities and is prompted to randomly substitute them, while the seek model is trained with the same data and the addition of the remote LLM answer. The framework is evaluated using both adversarial models (black box and white box) to measure privacy protection effectiveness and usability experiments to assess the impact on LLM task performance, including translation and classification. Adversarial models are unable to recover original information, indicating stronger privacy protection. Usability is evaluated by metrics like BLEU for translation and precision, recall, and F1 scores for classification, to see how well the framework maintains the utility of LLM outputs post-anonymization. The results demonstrate that HaS effectively balances privacy and utility, with label-based anonymization showing better privacy protection and generative methods offering slightly better utility retention.
@@ 4.1.2 Anonymization with Differential Privacy
In this section, we are proposing works that make use of Differential Privacy for data anonymization i.e., before actually using the data in training. We refer the reader to §2.1 for an introduction to Differential Privacy, and we will propose later in §5 methods that use Differential Privacy directly on the model. The methods proposed in this section apply Differential Privacy in the form of text rewriting: a text is modified to preserve privacy by perturbing the original text representations. The main advantage is that the Differential Privacy formal guarantees are met. All the works in this section are also available in tabular format in Table 3. Furthermore, referring to Figure 6, we identify the works in this section as examples of Differential Privacy applied on the input (1), as they all use DP to pre-process the input data before actually feeding it to the LLM, regardless of this being for training or inference.
@@
The Differential Privacy mechanism, in textual examples, should not disrupt the syntactic structure of a sentence. For this reason, Krishna et al. (2021) developed ADePT, an auto-encoder-based differential privacy algorithm aimed at anonymizing text while preserving its utility. This method involves clipping and noise introduction—either Laplacian or Gaussian—between encoding and decoding to form a differentially private latent representation. They hypothesize that data regeneration via an auto-regressive decoder maintains the syntactic structure of the sentence. Since they focus on the anonymization for Intention Classification (IC), an intention label is included before encoding to facilitate training an annotation-aware auto-encoder. Despite ADePT’s design to enhance privacy against Membership Inference Attacks and its superiority over traditional privacy-sensitive word replacement methods like Feyisetan et al. (2019b), Habernal (2021) point out that ADePT actually fails to achieve true differential privacy and exhibits higher sensitivity than reported, casting doubt on its effectiveness. In a similar fashion, Weggenmann et al. (2022) introduce a DP-VAE (Differentially Private Variational AutoEncoder) that also perturbs the latent space to secure privacy, following the model depicted in 5. This model contracts the mean vector of each latent representation towards the latent space’s origin before adding isotropic Gaussian noise. Applied to anonymize IMDb and Yelp reviews, this technique successfully reduces authorship attribution accuracy significantly while only slightly impacting sentiment classification accuracy. Further testing with a GPT2-based model confirms that the anonymized texts remain coherent and understandable, despite a slight increase in perplexity.
@@
Also transformer-based models can be used in this setting. Extending the research on Differential Privacy for text privatization in this direction, Igamberdiev & Habernal (2023) introduced DP-BART, a privatization model using the BART architecture instead of an LSTM, employing a similar strategy of clipping and noise addition in the latent encoder output, as seen in 5. DP-BART tackles the challenges of high sensitivity and computational demands associated with transformers through a novel clipping technique (by value) and an iterative pruning and training method that simplifies the encoder’s output vectors. Tested against ADePT on datasets like ATIS, Snips, IMDb, Drugs.com, and Amazon, DP-BART demonstrated superior performance in both privacy and utility. Another transformer-based approach is proposed by Du et al. (2023a), where they use BERT and propose the Purkayastha Mechanism to take advantage of the angular distance as a metric. They also propose to sanitize labels when too sensitive, using Randomized Response if the label space is small enough. Otherwise, they prune the label space leveraging prior knowledge and then use RR. This approach is tested on SST-2, IMDb, and QNLI datasets, representing tasks for sentiment analysis and question-answering based on movie reviews and Wikipedia content, respectively. Performances are close to the non-private version with an increased robustness to MIAs.
@@
Yue et al. (2021) argue that – due to the curse of dimensionality – directly injecting noise in large high-dimensional word embedding may fail to strike a nice privacy-utility balance, since it becomes exponentially less likely to find a noisy embedding close to a real one on every dimension. Yue et al. (2021) developed two strategies, SANTEXT and SANTEXT+, to replace original text tokens with "sanitized" tokens via the exponential mechanism. Both methods utilize embeddings to map words to compute their semantic distance. Then, once a word is identified to be replaced, the target word is selected with a probability that is inversely proportional to the distance with respect to the original one.SANTEXT indiscriminately replaces any word, often resulting in excessive utility loss and protection, while SANTEXT+ specifically targets and replaces the top w low-frequency tokens, identified as most likely to be private. These approaches enforce Metric Local Differential Privacy (MLDP), maintaining utility by favouring tokens closer to the original, though requiring a high epsilon value, like 16, to achieve notable improvements in defence rate. Similarly, Chen et al. (2023a) introduced CUSTEXT, another MLDP approach where every token is replaced using a mapping function that determines the output set for an input token and a sampling function based on the exponential mechanism for selecting replacements. The main difference here is that every input embedding has a custom set of output embeddings, which is a subset of the full outputs space used in SANTEXT and thus it reduces the chances that the output is semantically irrelevant to the input. They claim the best privacy/accuracy tradeoff for these systems so far, achieving good results with relatively low epsilon values (up to 3), although there is still a notable performance drop compared to using plain text.
@@
SANTEXT and CUSTEXT, while being effective, show some limitations. Tong et al. (2024) critique the small adjacency list used in SANTEXT and CUSTEXT, which could expose the methods to embedding inversion attacks and inefficacy in open-ended text generation, where bias matters more. Tong proposes RANTEXT, which dynamically generates a "random adjacency list" for each token then adds noise to the token’s embedding and selects nearby tokens based on the noisy embedding, leveraging the Laplace mechanism to ensure Differential Privacy. This method, developed within a framework for using black-box LLMs in private inference, includes a denoising module to reconstruct private outputs. RANTEXT, tested on CNN/Daily Mail (Hermann et al., 2015), Wikitext-103-v1 (Merity et al., 2016) and the ArXiv dataset (Clement et al., 2019), generally outperforms other methods in both utility and privacy, with exceptions in some areas where CUSTEXT+ performs slightly better. This comprehensive approach enhances the robustness of textual differential privacy, especially against inference and embedding inversion attacks.
@@
Expanding on the theme of embedding-based perturbation methods, Xu et al. (2021) acknowledge a common limitation in such approaches, where even minor noise might still leave the original embedding as the nearest neighbour. To counter this, they introduce a method inspired by Vickrey auctions to balance the selection between the first and second closest neighbours using a tuning parameter. This parameter is optimized through a constrained optimization problem aiming to maximize utility while ensuring privacy guarantees. Conducted experiments on Product Reviews, IMDb Movie Reviews, and Twitter datasets reveal that this Vickrey mechanism can enhance utility by up to 50% on real text classification datasets compared to existing methods, without compromising privacy, highlighting an innovative stride in balancing the privacy-utility tradeoff in textual differential privacy strategies. Further examining the methods in the field, Arnold et al. (2023) critique the focus on semantic-based embeddings in existing Metric Differential Privacy techniques, noting their failure to consider syntax specifically. Analyzing these techniques’ ability to preserve grammatical properties, they found that privatization mechanisms often privilege nouns even when replacing other grammatical categories. To address this, they incorporated grammatical categories into the privatization process, using methodologies from Feyisetan et al. (2019a) or Carvalho et al. (2021). Results indicate that this approach can enhance performance on downstream tasks, although it also shows a slight increase in self-substitution, suggesting a small decrease in privacy compared to previous approaches. This development underscores the complex interplay between maintaining linguistic integrity and ensuring robust privacy protection in textual data. Adding to the critique, Mattern et al. (2022b) identify significant limitations of Metric Differential Privacy methods, such as the increased noise necessary for long texts or their inability to rephrase effectively. Proposing an alternative, they suggest a paraphrasing model obtained by fine-tuning GPT2, inspired by Witteveen & Andrews (2019). This model introduces noise by sampling temperature based on the exponential mechanism, aiming to balance privacy with natural language generation capabilities. Tested for its effectiveness using the Matthews Correlation Coefficient (MCC) for authorship attribution and sentiment analysis, as well as semantic similarity and perplexity metrics to assess how well the original meaning is preserved and the quality of the generated text, this paraphrasing model proves superior to traditional word-level DP mechanisms.

### v2

@@ 4.1.1 Classical Anonymization
As already mentioned, anonymity definitions can be traced back to 1997 with k-anonymity. Anonymization aims to make the re-identification of an individual impossible, that is, the process of reconstructing the identity of an individual given a record that contains their information. Different attributes could be anonymized. While effectively masking PII – like names, emails, and social security numbers – is necessary, some features that are not exactly identifiers can be used in combination to identify an individual: those features are called quasi-identifiers. Quasi-identifiers are apparently non-personal attributes – like gender, date of birth, and zip codes – that can be used in combination with external resources to identify a person from its anonymized records. The definition of k-anonymity (Sweeney, 1997) ensures that a quasi-identifier cannot be used for re-identification. Essentially, a dataset is said to have k-anonymity if any given record in the dataset is indistinguishable from at least k - 1 other records with respect to certain identifying attributes called quasi-identifiers. Upon this concept, several implementation techniques were proposed by Samarati & Sweeney (1998), like generalization (modifying attribute values to less specific but accurate representations), suppression (i.e., entirely removing data that are outliers or too identifiable), minimal Generalization (ensuring data are not generalized more than necessary to meet k-anonymity). Of course, this method is highly susceptible to the choice of k (even inapplicable sometimes) and gives coarse data, making it impossible to carry out fine-grained analysis. It is also important to mention that, due to the reliance on quasi-identifiers, this method is much less effective with unstructured text.
@@
The standard process is outlined in Figure 4. After any necessary preprocessing phase, the first challenge to solve is to identify the entities that may cause a violation of privacy. This can be done via a Named Entity Recognition (NER) tool. To maintain the coherence of the text, a Co-reference Resolution (CRR) system can be used to resolve references between entities should be identified if the anonymization framework aims to ensure consistent anonymization across the document. Finally, private information should be replaced with anonymous information. The anonymization methods can vary. Suppression is used to completely remove the information. Tagging allows the replacement of sensitive information with artificial labels –like the one obtained by the NER system– that still retain information about the class of the data and an identifier. The sensitive information can also be replaced with other textual information: random substitution and generalization can both be used. Random substitution uses another random entity of the same class, while generalization replaces an entity with a more general term. It should be noted that, while this pipeline is intended for general data anonymization, regardless of which model is going to be trained on it (be that an LLM or not), it is specific to textual data, as it relies on some specific features of text.
Anonymization before DNNs
@@
While there is no doubt that these methods could be used to clean a dataset for LLM training or inference, it is not very clear how they would scale for an LLM-sized training dataset, as they are developed for much smaller datasets.
Anonymization with DNNs
More recent anonymization methods also leverage Deep Neural Networks. Zhou et al. (2022) use BERT to impute previously masked sensitive information. This framework is called DataSifterText, and it works with a whitelist that holds sensitive terms that need to be masked and a blacklist for meaningful terms that should be retained. After a proportion parameter is chosen, sensitive tokens are masked according to that proportion, and then BERT is used to replace the masked tokens. These two steps in the procedure respectively cover the identification and anonymization step shown in 4).
@@
In some cases, the author of a text is meant to be identified: re-identification of an author can be done by spotting some features that characterize an author’s writing. Romanov et al. (2019) analyze this scenario for the Russian language. They extract features like lemmas, parts of speech and morphological features. Then, a fast correlation filter – a multivariate feature selection method where the class relevance and the dependency between each feature pair are taken into account – is used to select the most informative features. Those informative features are the more likely to be identifying features. To preserve the privacy of the author, those features are smoothed by making targeted adjustments. For example, words or punctuation marks that are characteristic of the author’s style are replaced or altered based on average statistics derived from a broader corpus. The final step is to further process the smoothed text with a deep learning model: in this work, a Transformer-based model is chosen. This approach reduces the probability of authorship identification to random guessing.
@@
Mosallanezhad et al. (2019) use an attention-based task-aware text representation learner for extracting latent embedding representations with a reinforcement-learning-based privacy and utility preserver for manipulating these embeddings. While it effectively balances privacy and data utility and is adaptable and efficient, it has notable drawbacks, including complexity, potential data-specific limitations, and dependency on precise parameter tuning. Complexity in particular is relevant to this context, as it is likely to make this method less suitable for large-scale datasets.
@@
More recently, some works proposed the use of LLMs for anonymization. Vats et al. (2024) propose to hide sensitive tokens via a masking process and impute them by using a Large Language model (LLM). In relation to Figure 4, the masking acts as the identification part, as it finds and masks sensitive information, while the imputation represents the proper anonymization, where non-sensitive terms replace the (previously masked) sensitive ones. To carry out the masking process, several techniques are proposed: a handcrafted list of safe words (the ones that do not carry sensitive information) called allowlist created by linguistic experts, this is similar to the blacklist of Zhou et al. (2022) and identifies the terms that do not need to be masked; another method to identify safe words is vocabThres, where the N more frequent words in a vocabulary are considered safe; the last masking criteria is to find the unsafe words and involves using a NER module that identify potentially private tokens to mask them (similar to Mamede et al. (2016)). After choosing a masking method, the chosen tokens are masked and then imputed by an LLM. Also for the imputation, three separate methods are proposed: Top-1, which uses the one token with the highest probability from the LLM’s output; Top-K, which randomly selects one of the first K tokens with the highest output probabilities; Fine-Tuning, which selects the records that do not contain sensitive information and uses them to fine-tune the LLM, then just takes the highest probability output (so after fine-tuning it circles back to the Top-1 approach). They compare models trained on the anonymized datasets against models trained on the masked dataset, showing a significant improvement in both tasks. Intuitively, this method is quite sophisticated and complex, which means that for an LLM-sized training dataset it is going to make up for a quite heavy computation, but it could be suitable for a smaller dataset that might be used for fine-tuning or inference.
@@
As the usage of closed models increases, privacy concerns also emerge when sending prompts containing private information. Chen et al. (2023b) focus on sanitizing prompts before sending them out to some unsafe third party like a black-box model like ChatGPT, thus enforcing prompt privacy. Moreover, they argue that standard anonymization techniques – like masking – may be useful to ensure the privacy of prompts, but the model output could not be effectively de-anonymized to improve its utility. They propose a new framework called HaS (Hide and Seek) with two local small LMs, one trained for anonymization and one trained for deanonymization. The first one preprocesses the prompt before sending it to the black-box and the second one postprocesses the model output. This method is specifically designed to use LLMs in inference without worrying about personal data in the prompts. The hide model is trained with original sentences and extracted privacy entities and is prompted to substitute them randomly, while the seek model is trained with the same data and the addition of the remote LLM answer. The framework is evaluated using both adversarial models (black box and white box) to measure privacy protection effectiveness and usability experiments to assess the impact on LLM task performance, including translation and classification. Adversarial models are unable to recover original information, indicating stronger privacy protection. Usability is evaluated by metrics like BLEU for translation and precision, recall, and F1 scores for classification to see how well the framework maintains the utility of LLM outputs post-anonymization. The results demonstrate that HaS effectively balances privacy and utility, with label-based anonymization showing better privacy protection and generative methods offering slightly better utility retention.
@@ 4.1.2 Anonymization with Differential Privacy
In this section, we propose works that use Differential Privacy for data anonymization, i.e., before actually using the data in training. These methods are intended to anonymize datasets, regardless of the intended use of the datasets. This means that the anonymized datasets might be used to train LLMs, but also for other models and tasks. We refer the reader to §2.2 for an introduction to Differential Privacy, and we will propose later in §5 methods that use Differential Privacy directly on the model. The methods proposed in this section apply Differential Privacy to perform text rewriting, where the original text is modified by perturbing its representations to ensure privacy. The key advantage of this approach is that it adheres to the formal guarantees of Differential Privacy. All the works in this section are also available in tabular format in Table 3.
Referring to Figure 5, we identify the works in this section as examples of Differential Privacy applied to the input (1) of a neural model. Differential Privacy can also be applied in a separate, typically smaller, auxiliary preprocessing model to anonymize data before it is fed into an LLM.
Metric Differential Privacy
@@
The Differential Privacy mechanism, in textual examples, should not disrupt the syntactic structure of a sentence. For this reason, Krishna et al. (2021) developed ADePT, an autoencoder-based differential privacy algorithm aimed at anonymizing text while preserving its utility. This method involves clipping and noise introduction—either Laplacian or Gaussian—between encoding and decoding to form a differentially private latent representation. They hypothesize that data regeneration via an auto-regressive decoder maintains the syntactic structure of the sentence. Since they focus on the anonymization for Intention Classification (IC), an intention label is included before encoding to facilitate training an annotation-aware autoencoder. Despite ADePT’s design to enhance privacy against Membership Inference Attacks and its superiority over traditional privacy-sensitive word replacement methods like Feyisetan et al. (2019b), Habernal (2021) point out that ADePT actually fails to achieve true differential privacy and exhibits higher sensitivity than reported, casting doubt on its effectiveness. In a similar fashion, Weggenmann et al. (2022) introduce a DP-VAE (Differentially Private Variational AutoEncoder) that also perturbs the latent space to secure privacy, following the model depicted in 5. This model contracts the mean vector of each latent representation towards the latent space’s origin before adding isotropic Gaussian noise. Applied to anonymize IMDb and Yelp reviews, this technique successfully reduces authorship attribution accuracy significantly while only slightly impacting sentiment classification accuracy. Further testing with a GPT2-based model confirms that the anonymized texts remain coherent and understandable despite a slight increase in perplexity.
@@
Also transformer-based models can be used in this setting. Extending the research on Differential Privacy for text privatization in this direction, Igamberdiev & Habernal (2023) introduced DP-BART, a privatization model using the BART architecture instead of an LSTM, employing a similar strategy of clipping and noise addition in the latent encoder output, as seen in 5. DP-BART tackles the challenges of high sensitivity and computational demands associated with transformers through a novel clipping technique (by value) and an iterative pruning and training method that simplifies the encoder’s output vectors. Tested against ADePT on datasets like ATIS, Snips, IMDb, Drugs.com, and Amazon, DP-BART demonstrated superior performance in both privacy and utility. Another transformer-based approach is proposed by Du et al. (2023a), where they use BERT and propose the Purkayastha Mechanism to take advantage of the angular distance as a metric. They also propose to sanitize labels when too sensitive, using Randomized Response if the label space is small enough. Otherwise, they prune the label space, leveraging prior knowledge and then use RR. This approach is tested on SST-2, IMDb, and QNLI datasets, representing tasks for sentiment analysis and question-answering based on movie reviews and Wikipedia content, respectively. Performances are close to the non-private version with an increased robustness to MIAs.
Levaraging the Exponential Mechanism
@@
Yue et al. (2021) argue that – due to the curse of dimensionality – directly injecting noise in large high-dimensional word embedding may fail to strike a nice privacy-utility balance since it becomes exponentially less likely to find a noisy embedding close to a real one on every dimension. Yue et al. (2021) developed two strategies, SANTEXT and SANTEXT+, to replace original text tokens with "sanitized" tokens via the exponential mechanism. Both methods utilize embeddings to map words to compute their semantic distance. Then, once a word is identified to be replaced, the target word is selected with a probability that is inversely proportional to the distance with respect to the original one. SANTEXT indiscriminately replaces any word, often resulting in excessive utility loss and protection, while SANTEXT+ specifically targets and replaces the top w low-frequency tokens, identified as most likely to be private. These approaches enforce Metric Local Differential Privacy (MLDP), maintaining utility by favouring tokens closer to the original, though requiring a high epsilon value, like 16, to achieve notable improvements in defence rate. Similarly, Chen et al. (2023a) introduced CUSTEXT, another MLDP approach where every token is replaced using a mapping function that determines the output set for an input token and a sampling function based on the exponential mechanism for selecting replacements. The main difference here is that every input embedding has a custom set of output embeddings, which is a subset of the full output space used in SANTEXT, and thus, it reduces the chances that the output is semantically irrelevant to the input. They claim the best privacy/accuracy tradeoff for these systems so far, achieving good results with relatively low epsilon values (up to 3), although there is still a notable performance drop compared to using plain text.
@@
SANTEXT and CUSTEXT, while being effective, show some limitations. Tong et al. (2024) critique the small adjacency list used in SANTEXT and CUSTEXT, which could expose the methods to embedding inversion attacks and inefficacy in open-ended text generation, where bias matters more. Tong proposes RANTEXT, which dynamically generates a "random adjacency list" for each token, then adds noise to the token’s embedding and selects nearby tokens based on the noisy embedding, leveraging the Laplace mechanism to ensure Differential Privacy. This method, developed within a framework for using black-box LLMs in private inference, includes a denoising module to reconstruct private outputs. RANTEXT, tested on CNN/Daily Mail (Hermann et al., 2015), Wikitext-103-v1 (Merity et al., 2017) and the ArXiv dataset (Clement et al., 2019), generally outperforms other methods in both utility and privacy, with exceptions in some areas where CUSTEXT+ performs slightly better. This comprehensive approach enhances the robustness of textual differential privacy, especially against inference and embedding inversion attacks.
As already discussed, none of these methods is specific to LLMs but they all could be used to preprocess the training dataset for an LLM, although with some computational overhead given by the usual (huge) size of LLMs’ training datasets.
Beyond the exponential mechanism
@@
Known Issues
Expanding on the theme of embedding-based perturbation methods, Xu et al. (2021) acknowledges a common limitation in such approaches, where even minor noise could still leave the original embedding as the nearest neighbour. To counter this, they introduce a method inspired by Vickrey auctions to balance the selection between the first and second closest neighbours using a tuning parameter. This parameter is optimized through a constrained optimization problem aiming to maximize utility while ensuring privacy guarantees. Conducted experiments on Product Reviews, IMDb Movie Reviews, and Twitter datasets reveal that this Vickrey mechanism can enhance utility by up to 50% on real text classification datasets compared to existing methods without compromising privacy, highlighting an innovative stride in balancing the privacy-utility tradeoff in textual differential privacy strategies. Further examining the methods in the field, Arnold et al. (2023) critiques the focus on semantic-based embeddings in existing Metric Differential Privacy techniques, noting their failure to consider syntax specifically. Analyzing these techniques’ ability to preserve grammatical properties, they found that privatization mechanisms often privilege nouns even when replacing other grammatical categories. To address this, they incorporated grammatical categories into the privatization process, using methodologies from Feyisetan et al. (2019a) or Carvalho et al. (2021). Results indicate that this approach can enhance performance on downstream tasks, although it also shows a slight increase in self-substitution, suggesting a small decrease in privacy compared to previous approaches. This development underscores the complex interplay between maintaining linguistic integrity and ensuring robust privacy protection in textual data. Adding to the critique, Mattern et al. (2022b) identify significant limitations of Metric Differential Privacy methods, such as the increased noise necessary for long texts or their inability to rephrase effectively. Proposing an alternative, they suggest a paraphrasing model obtained by fine-tuning GPT2, inspired by Witteveen & Andrews (2019). This model introduces noise by sampling temperature based on the exponential mechanism, aiming to balance privacy with natural language generation capabilities. Tested for its effectiveness using the Matthews Correlation Coefficient (MCC) for authorship attribution and sentiment analysis, as well as semantic similarity and perplexity metrics to assess how well the original meaning is preserved and the quality of the generated text, this paraphrasing model proves superior to traditional word-level DP mechanisms. Although these methods show promise, it would be interesting to see more of the generated data used in tasks such as text generation, which is quite rare in the current literature. Most of the presented works focus on smaller tasks and models, but there is no limitation in using these methods to generate anonymized data to train LLMs.